a 20 kg box is initially at rest with a static and kinetic friction of 0.40 is pushed with a force of 70 For 120 s. What is the work done?


1
Expert's answer
2020-07-23T08:46:14-0400

"W=F\cdot s"


"F_{fr}=\mu mg=0.4\cdot 20\cdot 9.81=78.48N>70N"


This force ("70 N") will not be able to move the box. So,

"s=0\to W=0" Answer.